OpenStudy (xximhisgirlxx):

Suppose you live 10.2 miles from a hill. From your home you can see a plane directly above the hill. Your angle of elevation to the plane is 34 degrees. What is the planes altitude? 6.88 miles 8.46 miles 5.70 miles -0.23 miles

OpenStudy (xximhisgirlxx):

@mathstudent55 @thomaster

OpenStudy (anonymous):



OpenStudy (anonymous):

Use formula tan 34 =h/10.2 h=10.2*tan 34 h=6.88 miles ( rounding to 2 decimal places)

OpenStudy (anonymous):

Option a)